OpenStudy (anonymous):

The number of fish (f) in a fish tank increases when the number of algae blooms (a) decreases. Write the correct equation for this scenario, and solve for the number of fish when there are 4 algae blooms. algae blooms fish 10 10 25 4 a) f = 1a; f = 4 b) f = .1a; f = .4 c) f = 100 over a; f = 25 d) f = 100 over a; f = 400

OpenStudy (anonymous):

Your best bet is to substitute what you have into the chart to see if it works correctly with the equations.

OpenStudy (campbell_st):

well it seems like inverse variation \[f = \frac{k}{a}\] so substitute f 10 and a = 10 then solve for k. the variation constant. when you find k, substitute b = 4 to find the number of fish

OpenStudy (campbell_st):

oops substitute a = 4 to find the number of fish

OpenStudy (anonymous):

@campbell_st yeah i was like WAIT WHERE DID THE B COME FROM haha

OpenStudy (anonymous):

so it would be A

OpenStudy (campbell_st):

well what valeu did you get for k...?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

C seems to be the most reasonable one tbh. Cause 10 = 100/10 is correct and 4 = 100/25 is also correct

OpenStudy (anonymous):

im stupid

OpenStudy (campbell_st):

no... look at the equation again \[10 = \frac{k}{10}\] solve for k

OpenStudy (anonymous):

...100?

OpenStudy (anonymous):

dude 1

OpenStudy (anonymous):

its 1 isnt it

OpenStudy (campbell_st):

thats correct so the equation you will use is \[fish = \frac{100}{a}\] now substitute a = 4 to find the number of fish

OpenStudy (anonymous):

is a still 10 or is that gone now ?

OpenStudy (anonymous):

no 4

OpenStudy (anonymous):

25

OpenStudy (campbell_st):

no... you need to confirm the number of fish increases as the blooms decrease... to the other information you have been given is that the fish = 25 when the blooms = 4 so using the equation \[f = \frac{100}{a}\] substitute a = 4 and confirm f = 25

OpenStudy (anonymous):

but how is that confirmed ?

OpenStudy (campbell_st):

if you confirm its correct.... then find the equation in your answer choices and the fact that f = 25

OpenStudy (anonymous):

By solving the equation with a being 4

OpenStudy (anonymous):

ah so it is C ! like dariel said

OpenStudy (campbell_st):

substitution is the process of replacing a letter by a number so using a = 4 and substituting into \[f = \frac{100}{a}\] you will be able to calculate f \[f = \frac{100}{4}\] so what is f... if you do the calculation above...?

OpenStudy (anonymous):

Both our methods works, just mine is not always as reliable :3

OpenStudy (anonymous):

both fantastic helpers thanks ! ^.^

OpenStudy (anonymous):

no problem